# Impact of Transmission Dynamics and Treatment Uptake, Frequency and Timing on the Cost-effectiveness of Directly Acting Antivirals for Hepatitis C Virus Infection

### v2

@@ Abstract
Cost-effectiveness analyses, based on decision-analytic models of disease progression and treatment, are routinely used to assess the economic value of a new intervention and consequently inform reimbursement decisions for the intervention. Many decision-analytic models developed to assess the economic value of highly effective directly acting antiviral (DAA) treatments for the hepatitis C virus (HCV) infection do not incorporate the transmission dynamics of HCV, accounting for which is required to estimate the number of downstream infections prevented by curing an infection. In this study, we develop and validate a comprehensive agent-based simulation (ABS) model of HCV transmission dynamics in the Indian context and use it to: (a) quantify the extent to which the cost-effectiveness of a DAA is underestimated - as a function of its uptake rate - if disease transmission dynamics are not considered in a cost-effectiveness analysis model; and (b) quantify the impact of the frequency and timing of treatment within a disease surveillance period, also as a function of uptake rate, on DAA cost-effectiveness. The process of accomplishing the above research objectives also motivated the development of a novel random sampling and allocation approach, along with associated theoretical grounding, to estimate individual-level outcomes within an ABS that incurs substantially lower computational expense than the benchmark incremental accumulation approach.
@@ 1 Introduction
We accomplish the above research objectives by developing an agent-based simulation model (ABM) of HCV transmission dynamics in the Indian state of Punjab, which has one of the highest HCV prevalences in the country [Sood et al., 2018]. ABMs are a natural choice here [Ghaderi, 2022, Duma & Aringhieri, 2023, Nguyen et al., 2024] given their capability of incorporating the individual-based heterogeneity and stochasticity inherent to infectious disease transmission dynamics, and have previously been used to evaluate HCV treatment policy in the developed world [Ayer et al., 2019]. Using ABMs to evaluate the cost-effectiveness of interventions for diseases such as HCV requires simulating the progression of disease for each infected agent over their lifetime - in other words, lifetime horizon outcomes need to be estimated. This involves considerable computational expense given that (a) life expectancies are not affected substantially for most HCV patients (quality of life, however, is); (b) day-long time steps were found to be appropriate for advancing the ABM; (c) the run length of a single replication was 60 simulation years; and (d) a relatively large number of agents were considered in our ABM. Such computational expense can arise in ABMs of any infectious disease with natural histories significantly longer than the ABM or intervention evaluation time horizon; for instance, in ABMs developed to model HIV transmission or even tuberculosis (TB). This motivated our third (methodological) research objective ($RO3$): develop, with mathematical grounding, a computationally efficient method for estimating lifetime horizon outcomes, preserving their inherent stochasticity, for ABMs of infectious disease transmission dynamics.
@@
The remainder of the paper is organized as follows. In Section 2, we discuss the relevant literature and describe our research contributions. In Section 3, we describe the ABM development and our proposed method to reduce its computational expense (i.e., $RO3$). In Section 4, we discuss computational experiments conducted to address $RO1$ and $RO2$, and we conclude in Section 5 with a discussion of the potential impact of our study and avenues for future research.
@@ 2 Literature Review
In this section, we discuss the literature relevant to research objectives RO1 through RO3, and discuss our research contributions with respect to each objective. Prior to this, we provide a brief overview of the different types of models that have been developed for assessing the epidemiological, health and economic impact of HCV treatment and/or screening, with a focus on treatment with DAAs. Further, in Table 1 below, we describe the type of transmission model, transmission modes considered, type of disease progression model, the geography of analysis, whether single or multiple treatment uptake rates and frequencies were considered, whether timing of treatment was considered, the outcomes estimation approach, and the types of outcomes generated by the model. In other words, this table summarizes the literature relevant to all three research objectives.
@@ 2.2 Literature Pertaining to RO2: Impact of Intervention Frequency and Timing
2.3 Literature Pertaining to RO3: Outcomes Estimation in Agent-based Models
Two types of studies are commonly found in terms of the time horizon over which outcomes pertaining to cost-effectiveness analyses are evaluated [Basu & Maciejewski, 2019]. The first involves assessment of outcomes only within the duration of the intervention deployment period - these are called within-trial horizon analyses. In the context of an ABM, this would typically mean that outcomes for agents are assessed only until the end of the model execution time horizon. Alternatively, outcomes are assessed for a duration beyond the intervention deployment period. For ABMs, this often involves estimating outcomes for agents of interest for their lifetimes. Such analyses are called lifetime horizon analysis. We suggest a third type - fixed horizon analysis: one where outcomes are assessed beyond the end of the intervention deployment period and up to a certain pre-specified point of interest, but not for the entire duration of individual lifespans. Conducting lifetime horizon analyses with ABMs is challenging, as the execution of the disease progression model for agent lifespans for the entire agent cohort (typically large in size) may be computationally expensive or challenging to implement (the latter because termination points are likely to differ for each agent). Thus these studies are often within-trial or fixed horizon analyses. To address this, an outcomes allocation approach - e.g., assigning expected values of outcomes (e.g., quality-adjusted life years) generated from the DTMC model of the disease natural history - has also been used. However, this approach does not capture the stochasticity inherent in the progression of the disease. Implementing a stochasticity-preserving outcomes allocation approach within an ABM is a nontrivial task, as one has to consider how the sample path of an agent changes after an initial allocation - e.g., if an infected agent gets cured, and an initial allocation is made upon infection. If one allocates again at the point of change in the sample path (e.g., due to cure), how does one account for outcomes accumulated until this point? Discounting of outcomes also requires careful handling, as the reference point for discounting changes with each agent.
Many relevant studies (see Table 1) do not explicitly mention the exact method used for outcomes estimation, especially when a DTMC is used to simulate HCV natural history. In cases where they do, this approach is often an uncertainty-free one-time allocation of average outcomes generated by repeated multiplication of the transition probability matrix for the DTMC with the state distribution vector (starting from an initial distribution) for a fixed time period, as in Bennett et al. [2017]. Alternatively, outcomes estimation is done for fixed cohorts. We see similar approaches for outcomes estimation for agent-based models developed for other diseases as well, such as in HIV [Gopalappa et al., 2017] and HPV [Olsen & Jepsen, 2010].
In this study, as part of our research contribution associated with RO3, we propose a new and computationally more efficient approach, with theoretical grounding, for estimating the lifetime outcomes for agents of interest within an ABM that removes the need for executing the natural history DTMC (NH-DTMC) model beyond the end of the intervention deployment period. This approach takes into account changes in agent sample paths by integrating accumulated outcomes between points of change in agent sample paths with outcomes allocated at points of change. We call this approach the outcomes allocation (OA) method for lifetime horizon outcomes estimation.
@@ 3 Agent-Based Simulation Model Development
The ABM that we develop to accomplish research objectives 1 through 3 consists of the following components, each of which we describe in this section.
@@ 3.1 Agent Interaction Environments and Ancillary Dynamics
$q_{v}^{med}=\frac{m_{i}+m_{b}+m_{s}+m_{d}}{360}\$ (1)
In equation 1, $m_{j}$ denotes the annual average number of times the $j^{th}$ medical event of interest occurs, where $j$in {injections ($i$), blood transfusions   ($b$), surgeries ($s$), dental procedures ($d$)}. The proportion of medical procedures that were unsafe in the Indian context ($q_{uns}^{med}$) was estimated from Pala et al. [2016], which indicated that nearly half the clinics across the country did not follow safe sterilization practices, and thus $q_{uns}^{med}$ was set to 0.50. Next, the number of agents seen by a medical professional on a given day, denoted by $N_{p|mp}$, was estimated using the doctor-patient ratio ($r_{d|p}$), reported to be 1:1800 (see Table 2). We thus derived $N_{p|mp}$ using the relation $N_{p|mp}=\frac{q_{v}^{med}}{r_{d|p}}$. The number of medical professionals was increased with time in accordance with the increase in population.
@@ 3.2 Treatment Module for Infected Agents
$N_{j_{1},j_{2}}^{Tr}$ = Number of new infections caused at time point $j_{1}$ which are treated at time point $j_{2}$, where 0 $\leq$ $j_{1}$leq $j_{2}$leq $n$.
@@
In Algorithm 1, $j$ = 0 is the beginning of the period and $j$ = $n$ is the end of the intervention period, with treatment occurring over $n+1$ time periods. If $u$ is the target uptake rate across the intervention period, then we assume that a fraction $\alpha$ of $u$ (e.g., 30% of all those to be treated, which may be, say, 10% of all infected agents present across the intervention period) are treated at the end of $j=0$. The coverage of treatment in each time period is then scaled upwards linearly to reflect the likelihood of awareness increasing with time. In each time period $j>0$, we assign $u$of newly created infections to be treated ($N^{a}_{j})$, and then set the number treated in $j$ to the coverage fraction $\left[\alpha+(1-\alpha)\times\frac{j}{n}\right]$ multiplied by the total number of infections present at $j$, which is the sum of the number of infections remaining at the end of $j-1~{}(M^{l}_{j-1}$) and $N^{a}_{j}$. Note that if $\alpha\times u$ fraction of the infections present prior to $j=1$ are treated at $j=0$, then scaling up the coverage linearly implies that with each time period, a fraction of the remaining coverage (equal to $1-\alpha$) to be achieved must be added to the base coverage at $j=0$, $\alpha$. To ensure that coverage at $j=n$ is equal to $u$ (or in other words, $\alpha=1$), this fraction is set to $\frac{j}{n}\times(1-\alpha)$.
@@ 3.3 Outcomes Estimation Module
$t\leftarrow 0$triangleright $t$ is the point of call of this algorithm for an agent.
@@
     $a_{t}\leftarrow a_{t-1}$triangleright Update age group
     $Inf_{t},k_{t}\leftarrow\textit{Updt}(Inf,k)$triangleright Call Algorithm 4

### v3

@@ Abstract
Cost-effectiveness analyses, based on decision-analytic models of disease progression and treatment, are routinely used to assess the economic value of a new intervention and consequently inform reimbursement decisions for the intervention. Many decision-analytic models developed to assess the economic value of highly effective directly acting antiviral (DAA) treatments for the hepatitis C virus (HCV) infection do not incorporate the transmission dynamics of HCV, accounting for which is required to estimate the number of downstream infections prevented by curing an infection. In this study, we develop and validate a comprehensive agent-based simulation (ABS) model of HCV transmission dynamics in the Indian context and use it to: (a) quantify the extent to which the cost-effectiveness of a DAA is underestimated - as a function of its uptake rate - if disease transmission dynamics are not considered in a cost-effectiveness analysis model; and (b) quantify the impact of the frequency and timing of treatment within a disease surveillance period, also as a function of uptake rate, on DAA cost-effectiveness.
@@ 1 Introduction
The remainder of the paper is organized as follows. In Section 2, we discuss the relevant literature and describe our research contributions. In Section 3, we describe ABM development. In Section 4, we discuss computational experiments conducted to address $RO1$ and $RO2$, and we conclude in Section 5 with a discussion of the potential impact of our study and avenues for future research.
@@ 2 Literature Review
In this section, we discuss the literature relevant to research objectives RO1 and RO2, and discuss our research contributions with respect to each objective. Prior to this, we provide a brief overview of the different types of models that have been developed for assessing the epidemiological, health and economic impact of HCV treatment and/or screening, with a focus on treatment with DAAs. Further, in Table 1 below, we describe the type of transmission model, transmission modes considered, type of disease progression model, the geography of analysis, whether single or multiple treatment uptake rates and frequencies were considered, whether timing of treatment was considered, the outcomes estimation approach, and the types of outcomes generated by the model. In other words, this table summarizes the literature relevant to all three research objectives.
@@ 3 Agent-Based Simulation Model Development
The ABM that we develop to accomplish research objectives 1 and 2 consists of the following components, each of which we describe in this section.
@@ 3.1 Agent Interaction Environments and Ancillary Dynamics
$q_{v}^{med}=\frac{m_{i}+m_{b}+m_{s}+m_{d}}{360}\\$ (1)
In equation 1, $m_{j}$ denotes the annual average number of times the $j^{th}$ medical event of interest occurs, where $j$ $\in$ {injections ($i$), blood transfusions   ($b$), surgeries ($s$), dental procedures (dd)}. The proportion of medical procedures that were unsafe in the Indian context ($q_{uns}^{med}$) was estimated from Pala et al. [2016], which indicated that nearly half the clinics across the country did not follow safe sterilization practices, and thus $q_{uns}^{med}$ was set to 0.50. Next, the number of agents seen by a medical professional on a given day, denoted by $N_{p|mp}$, was estimated using the doctor-patient ratio ($r_{d|p}$), reported to be 1:1800 (see Table 2). We thus derived $N_{p|mp}$ using the relation $N_{p|mp}=\frac{q_{v}^{med}}{r_{d|p}}$. The number of medical professionals was increased with time in accordance with the increase in population.
@@ 3.2 Treatment Module for Infected Agents
$N_{j_{1},j_{2}}^{Tr}$ = Number of new infections caused at time point $j_{1}$ which are treated at time point $j_{2}$, where 0 $\leq$ $j_{1}$ $\leq$ $j_{2}$ $\leq$ $n$.
@@
In Algorithm 1, $j$ = 0 is the beginning of the period and $j$ = $n$ is the end of the intervention period, with treatment occurring over $n+1$ time periods. If $u$ is the target uptake rate across the intervention period, then we assume that a fraction $\alpha$ of $u$ (e.g., 30% of all those to be treated, which may be, say, 10% of all infected agents present across the intervention period) are treated at the end of $j=0$. The coverage of treatment in each time period is then scaled upwards linearly to reflect the likelihood of awareness increasing with time. In each time period $j>0$, we assign uu% of newly created infections to be treated ($N^{a}_{j})$, and then set the number treated in $j$ to the coverage fraction $\left[\alpha+(1-\alpha)\times\frac{j}{n}\right]$ multiplied by the total number of infections present at $j$, which is the sum of the number of infections remaining at the end of $j-1~(M^{l}_{j-1}$) and $N^{a}_{j}$. Note that if $\alpha\times u$ fraction of the infections present prior to $j=1$ are treated at $j=0$, then scaling up the coverage linearly implies that with each time period, a fraction of the remaining coverage (equal to $1-\alpha$) to be achieved must be added to the base coverage at $j=0$, $\alpha$. To ensure that coverage at $j=n$ is equal to $u$ (or in other words, $\alpha=1$), this fraction is set to $\frac{j}{n}\times(1-\alpha)$.
@@ 3.3 Outcomes Estimation Module
$\leftarrow 0 \triangleright tt$ is the point of call of this algorithm for an agent.
@@
  $a_{t}\leftarrow a_{t-1}$ $\triangleright$ Update age group
  $Inf_{t},k_{t}\leftarrow\textit{Updt}(Inf,k)$ $\triangleright$ Call Algorithm 3